OpenStudy (anonymous):

How much heat is required to raise the temperature of 80.0 g of water from its melting point to its boiling point? I have this equation set up right here m=80.0g, C=4.184J/(g⋅∘C), and ΔT=100∘C.

OpenStudy (baconboy73):

Now you need to plug it into the \[q = mC \Delta T\] equation and get your answer for q in joules.

OpenStudy (unklerhaukus):

\[q \ \ = (80.0\ \text g)\left(4.184\ \frac{\text J}{\text g\ °\text C}\right)(100\ °\text C)\\ \quad = (80.0\times4.184\times100) \frac{\cancel{\text g}\ \text J\ \cancel{°\text C}}{\cancel{\text g\ °\text C}}\\ \quad = (80.0\times4.184\times100)\ \text J\\ \quad =\]

OpenStudy (anonymous):

thanks